The Nisit Corporation is considering a project with a five-year life. The project requires $70,000 of fixed assets (initial installed cash outlay) that are classified as five-year property for MACRS. Variable costs equal 65 percent of sales, fixed costs are $14,000, and the tax rate is 36 percent. What is the operating cash flow for Year 4 given the following sales estimates and MACRS depreciation allowance percentages?

Year 1 2 3 4 5
Sales $24,000 $26,000 $28,000 $30,000 $32,000
MACRS rate 20.00 32.00 19.20 11.52 11.52

Also, please show how to entering numbers in financial calculator HP bII+
Answer #1

Sales for year 4 = $30,000

Variable costs for year 4 = $30,000 *65% = $19,500

Fixed costs = $14,000

Depreciation = Cost of fixed assets*MACRS rate for year 4 = $70,000 * 11.52% = $8,064

Operating income before taxes = Sales – Variable costs – Fixed costs – Depreciation

Operating income before taxes = $30,000 - $19,500 - $14,000 - $8,064 = -$11,564

Tax savings on operating income = $11,564 * 36% = $4,163.04

Operating income after tax savings = -$11,564+$4,163.04 = -$7,400.96

Operating cash flow = Operating income after tax savings + Depreciation

Operating cash flow for yea r4 = -$7,400.96+$8,064 = $663.04

This amount can be entered in financial calculator using CFj against year 4 cash flow as positive value.